SX Monthly Checkout

16:17

  • Powered on, ready to go.
  • New Hexapod positions: 2.99 3.29 -1.098 -107.6 -16.493
  • Note: Had to center the N30 mask.

17:00

  • Optical gain is much higher than it should be (0.5 instead of 0.3) and I'm getting odd NCPA coefficient amplitudes, making the NCPA applied images look worse than the ones with NCPA off.
  • Therefore, I'm leaving it off, after this quick check.
  • Bin 1, 1700 Hz, /local/aomeas/adsec_calib/CMD/disturb/dist_flao2_KL_v20_atm_s2.3_L040.0_v15.0_ovfreq1700.00_sd3892.fits
  • 20210127_171219 No LUCI frames
  • 20210127_172452 LUCI frames, NCPA off
  • 20210127_172702 NCPA On

17:46

  • Bin 2 pupils still need to be shifted.
  • Bin 3
  • 750Hz
  • /local/aomeas/adsec_calib/CMD/disturb/dist_flao2_KL_v20_atm_s2.0_L040.0_v15.0_ovfreq750.00_sd3892.fits
  • I think that somebody is walking around with a flashlight.
  • No, it's the lower floods. They are causing excess light on the detector.
  • New preset, lower floods off.
  • Shutter 49
  • mag 15.5
  • 20210127_180650

18:07

  • Going to do bin 4
  • Shutter 45
  • Magnitude 16.6
  • 428 Hz
  • /local/aomeas/adsec_calib/CMD/disturb/dist_flao2_KL_v20_atm_s2.0_L040.0_v15.0_ovfreq500.0000_sd3892.fits
  • 20210127_182417

18:25

  • We have a new bin2 pupil to try: 20210127_180255 (shifted by 0, -1)
  • It seems to work. I've changed both the ESM and the ACE tables.
  • dist_flao2_KL_v20_atm_s2.0_L040.0_v15.0_ovfreq1000.00_sd3892.fits
  • 904Hz
  • 13.4 mag
  • 20210127_184436

18:45

  • Going to try the ESM Reconstructor
  • Same lamp settings, same disturb as previous bin2
  • 400 Hz
  • 13.4mag
  • 20210127_185308

18:54

  • I'm going to test rotation at bin2
  • wobble_test_reverse.py
  • It still rails a little bit
  • Start: 20210127_185847
  • Finish:20210127_190233
  • New Pupils:20210127_190622 (shifted by 0, -2)
  • They behave as expected, loop closed properly in both ACE and TTM, there should be plenty of room in the camera lens stage now.
  • 440 to 0
  • Start: 20210127_192532
  • End: 20210127_193916
  • 0 to 440
  • Start: 20210127_194111
  • End: 20210127_195423

19:51

  • On the way back to 440, it railed in the other direction, badly.
  • I think that the pupil 20210127_180255 (shifted by 0, -1) is better. I'm going to revert and try the rotation again.
  • 440 to 0
  • Start: 20210127_200758
  • End: 20210127_202133
  • 0 to 440
  • Start: 20210127_203322
  • End: 20210127_204708

20:54

  • Going back to bin 1 to see if GOPT is still acting funny.
  • Looks like it's normal now, Optical gain is 0.3, for the same disturbance....
  • 20210127_210117 No LUCI frames
  • My network connection is bogging down. I'm stopping for now.

-- GregoryTaylor - 27 Jan 2021

Topic revision: r2 - 29 Jan 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback